OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3).

OpenStudy (anonymous):

parallel means same slope as y = 2x + 2 m = 2 Using point-slope y - 3 = 2(x - 5) Is this equation form okay?